OpenStudy (anonymous):

integrate 4x^2/e^-2x

OpenStudy (fibonaccichick666):

What ideas do you have?

OpenStudy (anonymous):

hold on its 4x^2/e^-2x

OpenStudy (fibonaccichick666):

@bloop it's \(e^{-2x}\)

OpenStudy (anonymous):

integration by parts, only im confused on the concept

OpenStudy (fibonaccichick666):

ok good

OpenStudy (fibonaccichick666):

so first we want to make our lives as easy as possible let's state what the definition/formula for integration by parts is

OpenStudy (anonymous):

integral udv=uv-integral vdu

OpenStudy (fibonaccichick666):

good, so we have this. Do you understand how I brought up the denominator?\[\int(4x^2e^{2x})dx\]

OpenStudy (anonymous):

yes i see how you got that

OpenStudy (fibonaccichick666):

ok so now let's pick our u and dv

OpenStudy (anonymous):

u=4x^2 and dv=e^2x

OpenStudy (fibonaccichick666):

alright so now we need du and v

OpenStudy (anonymous):

du=8x*dx

OpenStudy (anonymous):

im kinda confused on how to get the integral of e^2x

OpenStudy (fibonaccichick666):

what's the deriv of e^x?

OpenStudy (anonymous):

its just e^x right?

OpenStudy (fibonaccichick666):

yep so now we have a constant in front of our function, just like 4x^2, if you were to integrate that what would you do?

OpenStudy (anonymous):

well if i were integrating 4x^2 i would get 4x^3/3 would the same rule apply to e^2x?

OpenStudy (fibonaccichick666):

yup (minus the exponent rule)tell me what you get

OpenStudy (anonymous):

e^2x+1/2x+1

OpenStudy (fibonaccichick666):

not quite ok so let me explain it this way let u be a function of x e^u=u'e^u

OpenStudy (fibonaccichick666):

(Derivative)

OpenStudy (fibonaccichick666):

so if we let k be a constant. e^(kx)=ke^(kx) is the derivative, so what is the integral?

OpenStudy (anonymous):

2e^2x?

OpenStudy (fibonaccichick666):

close, but that would be the derivative

OpenStudy (anonymous):

e^2x/2

OpenStudy (fibonaccichick666):

yup so now let's put that into our eq and what do we get?

OpenStudy (anonymous):

=uv-int(vdu) =4x^2*e^2x/2-int(e^2x/2*8x

OpenStudy (fibonaccichick666):

wait why do you have division?

OpenStudy (anonymous):

thats v right?

OpenStudy (fibonaccichick666):

in the integral

OpenStudy (fibonaccichick666):

but yea that's v

OpenStudy (anonymous):

v is e^2x/2 thats the division inside the integral

OpenStudy (fibonaccichick666):

you've written \[=4x^2*e^2x/2-\int(e^2x/2*8x)dx \] ohk sorry hard to read like that

OpenStudy (fibonaccichick666):

if you throw a \[ and a \)<--that should be a bracket around it, it will write it all fancy

OpenStudy (fibonaccichick666):

so now let's look at that integral

OpenStudy (fibonaccichick666):

Somplify and pull out the constants to start

OpenStudy (fibonaccichick666):

simplify*

OpenStudy (anonymous):

\[2x^2*e^2x-int(4x*e^2x)\)

OpenStudy (anonymous):

lmao the bracket thing didnt work

OpenStudy (fibonaccichick666):

\[2x^2*e^2x-int(4x*e^2x)\]

OpenStudy (fibonaccichick666):

that parenthesis needed to be a bracket

OpenStudy (fibonaccichick666):

now pull out the constant

OpenStudy (anonymous):

ok final answer 2x^2*e^2x-4*x^2/2*e^2x/2

OpenStudy (fibonaccichick666):

not quite but close you have one more thing

OpenStudy (fibonaccichick666):

\[2x^2*e^{2x}-4*\int xe^{2x}/2 dx\] this is actually what you have, you can't integrate that the way you did guess what you get to do again???!!??

OpenStudy (fibonaccichick666):

and oops that /2 shouldnt be there

OpenStudy (anonymous):

wait im confused again..

OpenStudy (fibonaccichick666):

ok after removing that 4, you have \(2x^2e^{2x}−4\int xe^{2x} dx\)

OpenStudy (fibonaccichick666):

do you see why?

OpenStudy (anonymous):

yea the integral constant rule

OpenStudy (fibonaccichick666):

ok now that integral, you have to use integration by parts... again :/

OpenStudy (anonymous):

the constant gets kicked to the outside of the integral

OpenStudy (fibonaccichick666):

yup the constant does

OpenStudy (anonymous):

ewww whyyyyyyyyy

OpenStudy (fibonaccichick666):

because it is a product haha

OpenStudy (fibonaccichick666):

so what is your u? dv?

OpenStudy (anonymous):

so okay u=x dv=e^2x du=1 v=e^2x/2

OpenStudy (agent0smith):

You have to use int. by parts again because it's still just as un-integrateable as before. Often in these cases you need to integrate by parts twice, sometimes more :S

OpenStudy (fibonaccichick666):

Good so apply, le int by parts and then solve just remember that -4 will multiply the whole thing

OpenStudy (fibonaccichick666):

So true @agent0smith

OpenStudy (fibonaccichick666):

btw I can't get to my questions for some reason right now they won't load so if i'm not here that's why